OpenStudy (anonymous):

What is the reference angle of 9pi/4 and 3pi/7 What is the reference angle of 9pi/4 and 3pi/7 @Mathematics

OpenStudy (anonymous):

pi/4

OpenStudy (vishal_kothari):

Suppose you have an angle theta and want to find the reference angle. The first thing you should do is find an equivalent angle in the range 0 <= new theta < 2pi. This is easy to do... If theta >= 2pi, then keep subtracting 2pi from theta until it is within range (remember, 2pi is one full revolution on the circle). Likewise, if theta < 0, keep adding 2pi to it until it falls within range. Once you have your new value of theta, compute the reference angle as follows: If the new theta is in the 1st Quadrant, then ref angle = theta. If the new theta is in the 2nd Quadrant, then ref angle = pi - theta. If the new theta is in the 3rd Quadrant, then ref angle = theta - pi. If the new theta is in the 4th Quadrant, then ref angle = 2pi - theta. So let's do your problems... 9pi/4 = (9/4)pi = (2.25)pi > 2pi So subtract 2pi from it to obtain a coterminal angle. (9/4)pi - (8/4)pi = (1/4)pi = pi/4 That falls within range, so your new theta is pi/4. Pi/4 is in the 1st Quadrant, so the reference angle is pi/4. 3pi/7 is already within range. This angle is also in the 1st Quadrant (3/7 < 1/2, so 3pi/7 < pi/2), so the reference angle is 3pi/7

OpenStudy (vishal_kothari):

thanks...